A machine part rotates at an angular speed of ωi. Its speed is then increased to ωf at an angular acceleration of α. If both the initial and final angular speeds are now doubled and the angular acceleration remains the same, by what factor is the angular displacement changed?
1
Expert's answer
2020-05-14T09:06:57-0400

Given data

Initial angular speed is ωi

Final angular speed is ωf

Angular acceleration is α.

The expression for the angular displacement of machine is

"\Delta\theta=\frac{((\omega_f)^2-(\omega_i)^2)}{(2\alpha)}" ........(1)

When the both the initial and final angular speeds are now doubled that is ( ω'i =2ωi) and ( ω'f=2ωf), then the new angular displacement is

"\Delta\theta '=\frac{((\omega'_ f)^2-(\omega'_i)^2)}{(2\alpha)}"

"=(\frac{(2\omega_f)^2-(2\omega_i)^2)}{(2\alpha) } 
 =\frac{4((\omega _f)^2-(\omega_i)^2))}{(2\alpha)}" ........ (2)

From equations (1) and (2), we get

"\Delta\theta '=4 \Delta\theta"

Hence , the angular displacement is increased by the factor 4.